thrice the middle number of three consecutive numbers is 30712.What is the sum of the first and the third numbers
Answer by JulietG(1812)

What do we know from the problem?
3*b = 30712
b = a+1
b = c-1
.
Solve for B
30712/3 = B
This does not work. B is not a whole number. Can you please check the numbers and information given?